[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon May 29, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue May 30 04:53:13 UTC 2017


HSU Shift Log for Tue May 30 2017 00:53:11 GMT-0400 (EDT)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Halperin guidance does not appear to be working.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170530/9f2f02ff/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list